OpenStudy (anonymous):

A number decreased by negative three is equal to -40. Write an equation and find the number

OpenStudy (anonymous):

Follow and break down the sentence. A number = x is decreased by -3 (take away -3) x - (-3) is equal to -40 x - (-3)=-40 x-(-3) = x+3 x+3 = -40 -3 -3 ___________ x+0 = -43 or x = -43